# Multifidelity linear regression for scientific machine learning from scarce data

### v1

@@ 2.2 Multifidelity Monte Carlo
The work [50] also shows that the optimal choice of the coefficients $\alpha_{k}$ is given by $\alpha_{k}=\frac{\rho_{1k}\sigma_{1}}{\sigma_{k}}$. Both the optimal model allocation $m_{1},\ldots,m_{K}$, and the optimal $\alpha_{k}$ depend on second-order statistics of the models $\sigma_{k}$ and $\rho_{1k}$. In practice, since these statistics are generally not known, they are estimated from pilot samples in order to determine the model allocation and coefficients $\alpha_{k}$, leading to sub-optimal $\alpha_{k}$ and $m_{k}$ choices that nevertheless have led to orders-of-magnitude cost reductions in multiple real-world applications, including subsurface resource management [41], space telescope engineering [6], and topology and aircraft design optimization [10, 30].
@@ 3 Multifidelity linear regression
In this section, we introduce a multifidelity linear regression approach based on a multifidelity Monte Carlo estimate of the unknown expectation $C_{XY}=\mathbb{E}[XY]$ in (5) rather than the standard single-fidelity Monte Carlo estimate. Section 3.1 presents our multifidelity linear regression setting and introduces our multifidelity estimator for the regression coefficients $\beta$. Section 3.2 presents and discusses several strategies for choosing the model allocations $m_{k}$ and the control variate coefficients $\alpha_{k}$ in the multifidelity covariance estimator. Section 3.3 provides optimality analysis of some choices for the control variate coefficients.
@@ 3.1 Multifidelity estimators
This section introduces our multifidelity linear regression approach: we will introduce the estimator for the general case of $K$ models $f^{(1)},\ldots,f^{(K)}$ with decreasing costs $w_{1}>w_{2}>\cdots w_{K}$ and decreasing fidelity $\rho_{11}>\rho_{12}>\cdots\rho_{1K}$, as before. We assume that we have $m_{1}$ high-fidelity data pairs, $\{(x_{i},y_{i}^{(1)})\}_{i=1}^{m_{1}}$, and $m_{2}>m_{1}$ low-fidelity data, $\{(x_{i},y_{i}^{(2)})\}_{i=1}^{m_{2}}$, and so on, i.e., $m_{k}>m_{k-1}$ data $\{(x_{i},y_{i}^{(k)}\}_{i=1}^{m_{k}}$ where $y_{i}^{(k)}=f^{(k)}(z_{i})$ and $x_{i}=x(z_{i})$. The $z_{i}$ are i.i.d. drawn from $\pi$ and note that the input data sets are nested: that is, the first $m_{1}$ inputs are evaluated by all models, the first $m_{2}$ by models $2,\ldots,K$, and so on. We assume that $C_{XX}$ is available (nearly) exactly, as discussed in Section 2.1.
We now introduce two different multifidelity estimators for $C_{XY}$, which in turn define multifidelity estimators for the regression coefficients $\beta$. Let $X_{m_{k}}\in\mathbb{R}^{d\times m_{k}}$ denote the matrix whose columns consist of the first $m_{k}$ inputs, and let $Y_{m_{k}}^{(j)}\in\mathbb{R}^{m_{k}}$ for $j\leq k$ denote the vector whose elements are given by the first $m_{k}$ outputs of model $j$.
First, let $\alpha_{k}\in\mathbb{R}$ for $k=2,\ldots,K$, and define
$\hat{C}_{XY}^{\alpha}$ $=\frac{1}{m_{1}}X_{m_{1}}Y^{(1)}_{m_{1}}+\sum_{k=2}^{K}\alpha_{k}\left(\frac{1}{m_{k}}X_{m_{k}}Y^{(k)}_{m_{k}}-\frac{1}{m_{k-1}}X_{m_{k-1}}Y^{(k)}_{m_{k-1}}\right),$ (11)
which defines
$\hat{\beta}_{\text{MF}}^{\alpha}=(C_{XX})^{-1}\hat{C}_{XY}^{\alpha}.$ (12)
Second, let $A_{k}\in\mathbb{R}^{d\times d}$ for $k=2,\ldots,K$, and define
$\hat{C}_{XY}^{A}=\frac{1}{m_{1}}X_{m_{1}}Y^{(1)}_{m_{1}}+\sum_{k=2}^{K}A_{k}\left(\frac{1}{m_{k}}X_{m_{k}}Y^{(k)}_{m_{k}}-\frac{1}{m_{k-1}}X_{m_{k-1}}Y^{(k)}_{m_{k-1}}\right),$ (13)
which defines
$\hat{\beta}_{\text{MF}}^{A}=(C_{XX})^{-1}\hat{C}_{XY}^{A}.$ (14)
Note that eq. 11 can be viewed as a special case of eq. 13 where $A_{k}=\alpha_{k}I$. We now show (a) that the estimators $\hat{C}_{XY}^{\alpha}$ and $\hat{C}_{XY}^{A}$ are unbiased estimators for $C_{XY}$, (b) that the estimators $\hat{\beta}_{\text{MF}}^{\alpha}$ and $\hat{\beta}_{\text{MF}}^{A}$ are unbiased estimators for $\beta^{*}$ given by (2), and (c) that the multifidelity learned models $\hat{f}(z;\hat{\beta}_{\text{MF}}^{\alpha})$ and $\hat{f}(z;\hat{\beta}_{\text{MF}}^{A})$ issue unbiased predictions with respect to the theoretical optimal model $\hat{f}(z;\beta^{*})$.
@@
Unbiasedness of multifidelity linear regression via covariate estimation:
@@
    $\mathbb{E}[\hat{C}_{XY}^{\alpha}]=\mathbb{E}[\hat{C}_{XY}^{A}]=\mathbb{E}[XY^{(1)}]$,
@@
    $\mathbb{E}[\hat{\beta}_{\rm MF}^{\alpha}]=\mathbb{E}[\hat{\beta}_{\rm MF}^{A}]=\beta^{*}$, and
@@
    $\mathbb{E}[\hat{f}(z;\hat{\beta}_{\rm MF}^{\alpha})]=\mathbb{E}[\hat{f}(z;\hat{\beta}_{\rm MF}^{A})]=\hat{f}(z;\beta^{*})$.
@@
We prove all three statements for the ‘$A$’ version of the estimator; the proof for the ‘$\alpha$’ version follows because (12) is a special case of (14) as noted above. The first statement follows from the definition of $\hat{C}_{XY}^{\alpha}$ and the linearity of expectations:
$\mathbb{E}[\hat{C}_{XY}^{\alpha}]$ $=\mathbb{E}\left[\frac{1}{m_{1}}X_{m_{1}}Y^{(1)}_{m_{1}}\right]+\sum_{k=2}^{K}A_{k}\left(\mathbb{E}\left[\frac{1}{m_{k}}X_{m_{k}}Y^{(k)}_{m_{k}}\right]-\mathbb{E}\left[\frac{1}{m_{k-1}}X_{m_{k-1}}Y^{(k)}_{m_{k-1}}\right]\right)$
@@
The second statement now follows because $\hat{\beta}_{\rm MF}^{\alpha}$ defined in (12) is linear in the estimate $\hat{C}_{XY}^{\alpha}$. The third statement then follows from linearity of the model in $\hat{\beta}_{\rm MF}^{\alpha}$. ∎
Remark. We emphasize again that we consider settings in which $C_{XX}$ is available exactly. If $C_{XX}$ in eqs. 12 and 14 is instead estimated from samples, then the first statement in Theorem 3.1 remains true, but the second and third statements of Theorem 3.1 no longer hold exactly.
@@ 3.2 Choices of multifidelity linear regression hyperparameters
Both of the multifidelity estimators we have defined depend on several hyperparameters: the numbers of high- and low-fidelity samples $m_{k}$, as well as the control variate coefficients $\alpha_{k}\in\mathbb{R}$ or $A_{k}\in\mathbb{R}^{d\times d}$ for $k=1,\ldots,K$. This section presents several strategies for setting these hyperparameters and summarizes our recommendations: Section 3.2.1 discusses choices of control variate coefficients and Section 3.2.2 discusses the sample allocations. Detailed analytical justifications for our recommendations are deferred to Section 3.3.
@@ 3.2.1 Control variate coefficients
We consider first the estimator (12), where the control variate coefficient $\alpha$ is a scalar. As we described in Section 2, the work [50] proves that the choice
$\alpha_{k}=\frac{\rho_{1,k}\sigma_{1}}{\sigma_{k}}=\frac{\operatorname{\mathbb{C}ov}[f^{(1)}(Z),f^{(k)}(Z)]}{\operatorname{\mathbb{V}ar}[f^{(k)}(Z)]}$ (15)
minimizes the MSE of the scalar MFMC mean estimator (9). To adapt this from the MFMC mean estimation setting, where the goal is to estimate $\mathbb{E}[f^{(1)}(Z)]$, to our multifidelity linear regression context where the goal is estimation of $C_{XY}=\mathbb{E}[XY^{(1)}]=\mathbb{E}[X\,f^{(1)}(Z)]$, let $g^{(k)}(z)=x(z)\,f^{(k)}(z)$, and define
$C_{1k}=\operatorname{\mathbb{C}ov}[g^{(1)}(Z),g^{(k)}(Z)]),\quad C_{kk}=\operatorname{\mathbb{C}ov}[g^{(k)}(Z),g^{(k)}(Z)]).$ (16)
In Section 3.3, we will show (in Theorem 3.4) that the choice
$\alpha_{k}^{*}=\frac{\mathsf{Tr}(C_{1k})}{\mathsf{Tr}(C_{kk})}$ (17)
@@
minimizes the generalized variance (trace of the covariance) of the multifidelity estimator (12). We will show in Section 3.3 that this in turn minimizes an upper bound on the conditional variance of the learned model predictions, $\operatorname{\mathbb{V}ar}[\hat{f}(Z;\hat{\beta}_{\rm MF}^{\alpha})|Z=z]$.
We now consider the estimator (14), where the control variate coefficient $A$ is a matrix. Then, we will show in Section 3.3, that with $C_{1k}$ and $C_{kk}$ defined in (16), the choice
$A_{k}^{*}=C_{1k}C_{kk}^{-1}$ (18)
minimizes the generalized variance of (14). Theorem 3.8 in Section 3.3 additionally shows that this in turn minimizes both the variance of the learned model predictions, $\operatorname{\mathbb{V}ar}[\hat{f}(Z;\hat{\beta}_{\rm MF}^{A})|Z=z]$ and the generalized variance of the coefficients $\hat{\beta}_{\rm MF}^{A}$.
Although the control variate coefficients given by eqs. 17 and 18 are optimal in the sense that they minimize the generalized variance of the estimators (12) and (14), respectively, they depend on the second-order statistics $C_{1k}$ and $C_{kk}$ defined in (16). In general, $C_{1k}$ and $C_{kk}$ are unknown, so they must be estimated from samples in order to use eqs. 17 and 18. Because $C_{1k}$ and $C_{kk}$ in (16) are $d\times d$ covariance matrices, obtaining good estimates of these statistics generally will require many more samples than estimating the the scalar statistics $\rho_{1,k}$, $\sigma_{1}$, and $\sigma_{k}$ which define (15). Since our focus is on the setting in which computational resources are limited, we thus propose the heuristic strategy of using the $\alpha$ choice defined in (15) which minimizes the variance of the MFMC mean estimator. We will show in our numerical results that this yields practical gains despite being theoretically less optimal than the choices (17) and (18), and performs similarly to the theoretically optimal choices in practical settings where the statistics $C_{1k}$ and $C_{kk}$ must be estimated.
To summarize, we have three possible strategies for defining the multifidelity control variate coefficient:
  • MF-mean strategy (general recommendation): use the scalar $\alpha_{k}$ defined in (15) which was derived for MFMC estimation of means, based on $\rho_{1,k}$, $\sigma_{1}$, and $\sigma_{k}$ either known or estimated from samples,
  • MF-$\alpha^{*}$ strategy: use the scalar $\alpha$ defined in (17), based on $C_{1k}$ and $C_{kk}$ either known or estimated from samples,
  • MF-$A^{*}$ strategy (recommended when $C_{1k}$ and $C_{kk}$ are known exactly): use the matrix $A_{k}$ defined in (18), based on $C_{1k}$ and $C_{kk}$ either known or estimated from samples.
In the rare cases where $C_{1k}$ and $C_{kk}$ are known, we recommend the MF-$A^{*}$ strategy because it yields an estimator (14) with lower generalized variance than the MF-$\alpha^{*}$ strategy for (12), which in turn has lower generalized variance than the MF-mean strategy for (12). In our numerical results, we will show that the MF-$A^{*}$ strategy is indeed optimal when the statistics $C_{1k}$ and $C_{kk}$ are known or can be estimated very accurately, but that this ‘optimal’ strategy can perform badly when these statistics must be estimated from a limited number of samples. Our numerical results will show that the MF-mean strategy typically yields multifidelity linear regression estimators with orders-of-magnitude variance reduction, despite being sub-optimal. Because the MF-mean strategy requires the estimation of simpler model statistics than are required for the two ‘optimal’ strategies, it is our general recommendation.
@@ 3.2.2 Sample allocation
There are two scenarios in which our multifidelity linear regression approach can be employed: in the first, the $m_{1}$ high-fidelity data and $m_{k}$ low-fidelity data for $k=2,\ldots,K$ are fixed and given, for example if studies of the model output’s dependence on the input variables have already been run and their results saved. In this case, we then choose one of the $\alpha_{k}$ or $A_{k}$ choices discussed in Section 3.2.1 and implement the multifidelity $\beta$ estimator with the given sample allocation $m_{k}$.
In the second scenario, the user starts with no data but can choose to run the models $f^{(1)},\ldots,f^{(K)}$ at different inputs to generate output data $y_{i}^{(k)}$, subject to constraints on the overall cost of generating the output data. The allocation given by eq. 10 for MFMC mean estimation from [50] minimizes the mean squared error of the MFMC estimator. This raises the question of how to adapt the optimal sample allocation approach from MFMC mean estimation to multifidelity linear regression. A natural way to do this is to minimize the generalized variance of the vector multifidelity estimator for $C_{XY}\in\mathbb{R}^{d}$. As we will show in Section 3.3, this would also minimize the variance of the learned model predictions, conditioned on the input variable $z$. However, just as the MSE-minimizing sample allocation for MFMC mean estimation in (10) depends on second-order statistics of the scalars $Y^{(k)}=f^{(k)}(Z)$ which must often be estimated in practice, any attempt to minimize the generalized variance of the vector MFMC estimators (12) and (14) will depend on second-order statistics of the vectors $XY^{(k)}$ which must be estimated. Unfortunately, these second-order statistics are then given by $d\times d$ covariance matrices, which require more samples to estimate well than the second-order scalar statistics required for eq. 10. Because poor estimates of these statistics will lead to sub-optimal sample allocations anyway, instead of developing an analogue of (10) for the multifidelity linear regression problem, we instead propose the heuristic strategy of using (10) as-is to determine the sample allocation for multifidelity linear regression. We will show in our numerical results that despite not being tailored to the multifidelity linear regression problem, the strategy (10) leads to significantly more robust and accurate models learned from scarce data than standard models trained from high-fidelity only.
To summarize, there are two different ways we recommend choosing the sample allocation $m_{k}$ in the estimator definitions in eqs. 12 and 14, depending on the usage scenario:

### v2

@@ 2.2 Multifidelity Control Variate Estimators
Both the optimal sample allocation $m_{1},\ldots,m_{K}$, and the optimal $\alpha_{k}$ depend on second-order statistics of the models $\sigma_{k}$ and $\rho_{1k}$. In practice, since these statistics are generally not known, they are estimated from pilot samples in order to determine the model allocation and coefficients $\alpha_{k}$, leading to sub-optimal $\alpha_{k}$ and $m_{k}$ choices that nevertheless have led to orders-of-magnitude cost reductions in multiple real-world applications, including subsurface resource management [48], space telescope engineering [7], and topology and aircraft design optimization [11, 37]. In these applications, low-fidelity models with correlation as low as 38% have been successfully used to reduce the variance of Monte Carlo estimators [7].
One key difference between MFMC and the more general ACV and MLBLUE estimation strategies is in the sampling strategy: note that the MFMC estimator (9) has a nested sampling strategy where the first $m_{k}$ inputs are evaluated at all models $f^{(1)},\ldots,f^{(k)}$. The ACV and MLBLUE estimators remove this constraint on the sampling strategy, and can achieve greater variance reductions than MFMC as a result. However, the optimal sampling strategy for both ACV and MLBLUE estimators still depends on second-order statistics for the models, and generally does not have an analytical solution and is instead obtained through numerical optimization methods, see [27, 65, 13] for details.
@@ 3 Multifidelity linear regression
In this section, we introduce a multifidelity linear regression approach based on using multifidelity control variate estimators for the linear regression estimator (5) rather than the standard high-fidelity only Monte Carlo estimate. Section 3.1 presents our multifidelity linear regression setting and introduces our multifidelity linear regression strategy. Section 3.2 presents and discusses several strategies for choosing the model allocations $m_{k}$ and the control variate coefficients $\alpha_{k}$ in the multifidelity covariance estimator. Section 3.3 provides optimality analysis of some choices for the control variate coefficients.
@@ 3.1 Multifidelity estimators
This section introduces our multifidelity linear regression approach: we will introduce the estimator for the MFMC case of $K$ models $f^{(1)},\ldots,f^{(K)}$ with decreasing costs $w_{1}>w_{2}>\cdots w_{K}$ and decreasing fidelity $\rho_{11}>\rho_{12}>\cdots\rho_{1K}$, as before. We again emphasize that we use the MFMC estimator here to simplify the exposition, and note that our main idea applies equally well to other multifidelity control variate estimators like ACV and MLBLUE [27, 65].
We assume that we have $m_{1}$ high-fidelity data pairs, $\{(x_{i},y_{i}^{(1)})\}_{i=1}^{m_{1}}$, and $m_{2}>m_{1}$ low-fidelity data, $\{(x_{i},y_{i}^{(2)})\}_{i=1}^{m_{2}}$, and so on, i.e., $m_{k}>m_{k-1}$ data $\{(x_{i},y_{i}^{(k)}\}_{i=1}^{m_{k}}$ where $y_{i}^{(k)}=f^{(k)}(z_{i})$ and $x_{i}=x(z_{i})$. We assume that $C_{XX}$ is available (nearly) exactly, as discussed in Section 2.1. Let $X_{m_{k}}\in\mathbb{R}^{d\times m_{k}}$ denote the matrix whose columns consist of the first $m_{k}$ inputs, and let $Y_{m_{k}}^{(j)}\in\mathbb{R}^{m_{k}}$ for $j\leq k$ denote the vector whose elements are given by the first $m_{k}$ outputs of model $j$. Let $A_{k}\in\mathbb{R}^{d\times d}$ for $k=2,\ldots,K$, and define the following multifidelity estimate of $c_{XY}$:
$\hat{c}_{XY}^{\rm MF}=\frac{1}{m_{1}}X_{m_{1}}Y^{(1)}_{m_{1}}+\sum_{k=2}^{K}A_{k}\left(\frac{1}{m_{k}}X_{m_{k}}Y^{(k)}_{m_{k}}-\frac{1}{m_{k-1}}X_{m_{k-1}}Y^{(k)}_{m_{k-1}}\right),$ (11)
which defines a multifidelity estimate for the linear regression model:
$\hat{f}(z;\hat{\beta}^{\rm MF})=x(z)^{\top}\hat{\beta}^{\rm MF},\qquad\text{where }\hat{\beta}^{\text{MF}}=(C_{XX})^{-1}\hat{c}_{XY}^{\rm MF}.$ (12)
We note that it is equally possible to directly define a multifidelity control variate estimator for the unknown model parameters $\hat{\beta}$ or for the learned model $\hat{f}(z;\hat{\beta})$ itself: these estimators are closely related to the approach defined in eqs. 11 and 12, which we discuss in more detail in Appendix A. For now, we focus solely on the estimators defined by eqs. 11 and 12.
We now show (a) that the estimator $\hat{c}_{XY}^{\rm MF}$ is an unbiased estimator for $C_{XY}$, (b) that the estimator $\hat{\beta}_{\text{MF}}$ is an unbiased estimator for $\beta^{*}$, and (c) that the multifidelity learned models $\hat{f}(z;\hat{\beta}_{\text{MF}})$ issues unbiased predictions with respect to the theoretical optimal model $\hat{f}(z;\beta^{*})$.
@@
Unbiasedness of multifidelity linear regression approach:
@@
    $\mathbb{E}[\hat{c}_{XY}^{\rm MF}]=\mathbb{E}[XY^{(1)}]$,
@@
    $\mathbb{E}[\hat{\beta}^{\rm MF}]=\beta^{*}$, and
@@
    $\mathbb{E}[\hat{f}(z;\hat{\beta}^{\rm MF})]=\hat{f}(z;\beta^{*})$.
@@
The first statement follows from the definition of $\hat{c}_{XY}^{\rm MF}$ and the linearity of expectations:
$\mathbb{E}[\hat{c}_{XY}^{\rm MF}]$ $=\mathbb{E}\left[\frac{1}{m_{1}}X_{m_{1}}Y^{(1)}_{m_{1}}\right]+\sum_{k=2}^{K}A_{k}\left(\mathbb{E}\left[\frac{1}{m_{k}}X_{m_{k}}Y^{(k)}_{m_{k}}\right]-\mathbb{E}\left[\frac{1}{m_{k-1}}X_{m_{k-1}}Y^{(k)}_{m_{k-1}}\right]\right)$
@@
The second statement now follows because $\hat{\beta}^{\rm MF}$ defined in (12) is linear in the estimate $\hat{c}_{XY}^{\rm MF}$. The third statement then follows from linearity of the model in $\hat{\beta}^{\rm MF}$. ∎
Remark. We emphasize again that we consider settings in which $C_{XX}$ is available exactly. If $C_{XX}$ in eq. 12 is instead estimated from samples, then the first statement in Theorem 3.1 remains true, but the second and third statements of Theorem 3.1 no longer hold exactly.
@@ 3.2 Choices of multifidelity linear regression hyperparameters
We note that our multifidelity regression approach depends on several hyperparameters: the numbers of high- and low-fidelity samples $m_{k}$, as well as the control variate coefficients $A_{k}\in\mathbb{R}^{d\times d}$ for $k=1,\ldots,K$. This section presents several strategies for setting these hyperparameters and summarizes our recommendations: Section 3.2.1 discusses choices of control variate coefficients and Section 3.2.2 discusses the sample allocations for the MFMC-type estimator introduced in Section 3.1. Detailed analytical justifications for our recommendations are deferred to Section 3.3. Section 3.2.3 briefly discusses choices of hyperparameters in alternative multifidelity control variate estimation strategies.
@@ 3.2.1 Control variate coefficients
We consider three different choices for the control variate coefficient $A_{k}$ in (11). First, we propose the following heuristic choice:
[…]
As we described in Section 2, the work [57] proves that $\alpha_{k}^{\rm mean}$ minimizes the MSE of the scalar MFMC mean estimator (9) — but this choice is not optimal for minimizing the variance of our $\hat{c}_{XY}$ estimator. Despite its sub-optimality, our numerical results show that this heuristic strategy nevertheless yields practical gains.
To define an optimal choice of $A_{k}$ for our multifidelity linear regression estimators, note that our goal is to estimate $C_{XY}=\mathbb{E}[XY^{(1)}]=\mathbb{E}[X\,f^{(1)}(Z)]$. Let $g^{(k)}(z)=x(z)\,f^{(k)}(z)$, and define
$\Gamma_{1k}=\operatorname{\mathbb{C}ov}[g^{(1)}(Z),g^{(k)}(Z)]),\quad\Gamma_{kk}=\operatorname{\mathbb{C}ov}[g^{(k)}(Z),g^{(k)}(Z)]).$ (14)
In Section 3.3, we will show (Theorem 3.4) that the choice
[…]
minimizes the generalized variance of (12). Theorem 3.4 in Section 3.3 additionally shows that this in turn minimizes both the variance of the learned model predictions, $\operatorname{\mathbb{V}ar}[\hat{f}(Z;\hat{\beta}^{\rm MF})|Z=z]$ and the generalized variance of the coefficients $\hat{\beta}^{\rm MF}$. However, we note that the optimal choice (15) depends on the second-order statistics $\Gamma_{1k}$ and $\Gamma_{kk}$, which generally are unknown and must be estimated from samples. Because $\Gamma_{1k}$ and $\Gamma_{kk}$ in (14) are $d\times d$ covariance matrices, obtaining good estimates of these statistics generally will require many more pilot samples than estimating the scalar statistics $\rho_{1,k}$, $\sigma_{1}$, and $\sigma_{k}$ which define (13). For this reason, we also consider restricting $A_{k}$ to be a scalar multiple of the identity: $A_{k}=\alpha_{k}I$ (similar to the heuristic choice), i.e., essentially using a scalar coefficient $\alpha_{k}\in\mathbb{R}$ instead of the matrix coefficient $A_{k}\in\mathbb{R}^{d\times d}$. We will show in Theorem 3.6 that the choice
[…]
@@
minimizes the generalized variance (trace of the covariance) of the multifidelity estimator over all possible scalar coefficients $\alpha_{k}$. We will show in Section 3.3 that this in turn minimizes an upper bound on the conditional variance of the learned model predictions, $\operatorname{\mathbb{V}ar}[\hat{f}(Z;\hat{\beta}^{\rm MF})|Z=z]$. Note that (16) depends only on the traces of $\Gamma_{1k}$ and $\Gamma_{kk}$, which are scalar quantities which may be easier to estimate from pilot samples than the full covariance matrices $\Gamma_{1k}$ and $\Gamma_{kk}$ (and our numerical results in Section 4 support this claim).
We will compare the three choices for $A_{k}$ described above in our numerical experiments in Section 4 for both the case where model statistics are known and the case when they must be estimated from pilot samples. Our experiments will illustrate that the optimal matrix choice (15) leads to the lowest variance when the statistics $\Gamma_{1k}$ and $\Gamma_{kk}$ are known, but that this choice can perform poorly in settings when these covariances must be estimated, which is typically the case in practical applications. The optimal scalar choice (16) is more robust to sample estimates of $\Gamma_{1k}$ and $\Gamma_{kk}$ in our experiments, and is our general recommendation. However, our numerics also show that the sub-optimal heuristic choice (13) still consistently leads to significant variance reduction relative to the high-fidelity only case, and this choice may be preferred when the model statistics $\rho_{1k}$ and $\sigma_{k}$ are known or can be estimated more easily than $\mathsf{Tr}(\Gamma_{1k})$ and $\mathsf{Tr}(\Gamma_{kk})$.
@@ 3.2.2 Sample allocation
There are two scenarios in which our multifidelity linear regression approach can be employed: in the first, the $m_{1}$ high-fidelity data and $m_{k}$ low-fidelity data for $k=2,\ldots,K$ are fixed and given, for example if studies of the model output’s dependence on the input variables have already been run and their results saved. In this case, we then choose one of the $A_{k}$ choices discussed in Section 3.2.1 and implement the multifidelity $\beta$ estimator with the given sample allocation $m_{k}$.
In the second scenario, the user starts with no data but can choose to run the models $f^{(1)},\ldots,f^{(K)}$ at different inputs to generate output data $y_{i}^{(k)}$, subject to constraints on the overall cost of generating the output data. The allocation given by eq. 10 for MFMC mean estimation from [57] minimizes the mean squared error of the MFMC estimator. This raises the question of how to adapt the optimal sample allocation approach from MFMC mean estimation to multifidelity linear regression. A natural way to do this is to minimize the generalized variance of the vector multifidelity estimator for $c_{XY}\in\mathbb{R}^{d}$. As we will show in Section 3.3, this would also minimize the variance of the learned model predictions, conditioned on the input variable $z$. However, just as the MSE-minimizing sample allocation for MFMC mean estimation in (10) depends on second-order statistics of the scalars $Y^{(k)}=f^{(k)}(Z)$ which must often be estimated in practice, any attempt to minimize the generalized variance of the vector MFMC estimator (12) will depend on second-order statistics of the vectors $XY^{(k)}$ which must be estimated. Unfortunately, these second-order statistics are then given by $d\times d$ covariance matrices, which require more samples to estimate well than the second-order scalar statistics required for eq. 10. Because poor estimates of these statistics will lead to sub-optimal sample allocations anyway, instead of developing an analogue of (10) for the multifidelity linear regression problem, we instead propose the heuristic strategy of using (10) as-is to determine the sample allocation for multifidelity linear regression. We will show in our numerical results that despite not being tailored to the multifidelity linear regression problem, the strategy (10) leads to significantly more robust and accurate models learned from scarce data than standard models trained from high-fidelity only.
To summarize, there are two different ways we recommend choosing the sample allocation $m_{k}$ in the estimator definitions in eq. 12, depending on the usage scenario:
@@
3.2.3 Sample allocation and coefficient selection for other multifidelity control variate estimators
[…]